What is the solution?
3x - 2y = 6 and 5x + 3y = -9
1
Expert's answer
2013-02-19T07:50:28-0500
3x - 2y = 6 and 5x + 3y = -9
We can multiply the first equation by 5 and the second by 3. Thus, we obtain

15x-10y=30 and 15x+9y=-27
15x=10y+30 and 15x=-27-9yEquating right parts of equations above we get:

10y+30=-27-9y
19y=-57
Therefore y=-3
As we know 3x-2y=6
3x=2y+6
x=(2y+6)/3=(2*(-3)+6)/3=0
Thus, solution is x=0, y=-3